Let f be of bounded variation on [a,b] and define V is,

V(x)= {0 if x=a

V1(a,x) if a less than or equal to x.less than or equal to b

Then show that a point of continuity of f is also a point of continuity of V and conversely


1
Expert's answer
2021-03-16T08:13:36-0400

The total variation of α on the interval [a, b]:


"V_f(a,b)=sup\{\displaystyle\sum^n_{i=1}|f(x_i)-f(x_{i-1})|"

where "sup" is taken over all possible partitions of [a, b].

If "x" is a point of continuity of f, then there exists

"V_f(a,x)=sup\{\displaystyle\sum^n_{i=1}|f(x_i)-f(x_{i-1})|"

where "f(x_n)=f(x)". (Example: "V_f(a,x)=f(x)-f(a)" )

That is, "V_f(a,x)" is continious at point "x" .


If "x" is a point of continuity of "V_f(a,x)" , then "f(x_n)=f(x)" has definite value, so "f(x)" is continious at point "x" .